We are experiencing a loss of signal on our Dark fiber between Baltimore and Washington.
We have contacted our provider and ask for them to investigate and correct the situation.
Date: 2019-03-07 22:21:51 UTC Work is finished
Date: 2019-03-07 21:08:35 UTC Our fiber is now repaired and services are restored.
We are waiting confirmation that the work is finished.
Date: 2019-03-07 15:39:14 UTC The emergency maintenance has been complete.
OSP crews will now cut the cable and being working on repairs at that time.
Date: 2019-03-07 13:44:29 UTC OSP techs are finishing up the final migrations now. All remaining circuits will have to be repaired for services to restore. Due to the large impact of the cable it may take longer that 07:30 CST to process the emergency maintenance . We will advise as soon as the maintenance has been complete. More updates to follow.
Date: 2019-03-07 08:35:44 UTC crews remain onsite at this time working to get all services roll to new fiber to restore services
Date: 2019-03-07 05:49:52 UTC Fibers between Ashburn and Newark are repaired.
Technicians are still working on Ashburn Beauharnois
Date: 2019-03-07 05:08:56 UTC Current Update: splicing techs are still in the area continuing to work. They are not hands off yet.
Date: 2019-03-07 02:44:46 UTC Our Provider OSP tech are continuing to perform OTDR testing in an effort to locate damage on this cable. As a temporary fix they are looking into a possible reroute to a different cable.
Date: 2019-03-07 01:22:57 UTC Our Provider OSP tech are doing OTDR traces to pin-point the location.
Date: 2019-03-06 20:50:11 UTC Moreover we isolated all the links between Ashburn and Newark, the traffic to Ashburn is currently going through Chicago
Date: 2019-03-06 20:23:24 UTC we isolated Cogent, Equinix and Telefonica to avoid any saturation.